# Numerical efficiency of explicit time integrators for phase-field models

### v1

@@ 5.3 Grain growth with particles
The simulation with mobile particles show the triple lines (dark) defined as the 0.5 isolevel of the $\phi_{tj}=\min(27\sum_{i,j>i,k>j}\phi_{i}\phi_{j}\phi_{k},1)$ field together with the 0.5 isolevel of the particle phase (bluish transparent). The pores co-move together with the grain boundaries and merge over time, reducing their number. As indicated by the closeup views, pores can attain a variety of shapes depending on their environment and time to equilibrate. If the pores only touch two grains, their shape tends to be a squashed sphere, though most pores touch more grains. In that case the shape is close to a rounded polyhedron with vertices located at the enclosing grain triple lines in order to satisfy the surface tension balance; a tetrahedral example is shown for the middle row. This is in line with Tikare et al. [2001], HÖTZER et al. [2016], who employed a more regular geometry with pores being placed on quadruple points between a basal hexagonal polycrystal and one top crystal growing into this polycrystal. They observed that the pores took on shapes similar to ice-cream cones, with a triangular cross-section, which is the cross-section of a tetrahedron. Pores merging results in temporary irregular shapes as shown on the bottom. This pore is the result of the tetrahedral pore (four-sided face showing) from the middle row merging with the spheroidal pore shown in the middle panel and one nearby other tetrahedral pore. A video of the time evolution of this subdomain is deposited with the supplementary material. Due to the intentionally large pore mobility no pores are left behind within the growing grains.
@@
6 Conclusion and outlook
@@
What bears further investigation is the relatively large number of step rejections for some benchmarks. This suggests that either the stability limit is passed or that some part (error estimator, time step adapter) in the step adaption machinery fails. The former might be probed by including a dynamical estimate of the largest eigenvalue of the system. If this doesn’t affect the rejection rate significantly, it might be that standard adapters need significant tuning in their coefficients to reduce their rejection rate as shown by Ranocha et al. [2022] for compressible CFD problems.

### v2

@@ 5.3 Grain growth with particles
The simulation with mobile particles show the triple lines (dark) defined as the 0.5 isolevel of the $\phi_{tj}=\min(27\sum_{i,j>i,k>j}\phi_{i}\phi_{j}\phi_{k},1)$ field together with the 0.5 isolevel of the particle phase (bluish transparent). The pores co-move together with the grain boundaries and merge over time, reducing their number. As indicated by the closeup views, pores can attain a variety of shapes depending on their environment and time to equilibrate. If the pores only touch two grains, their shape tends to be a squashed sphere, though most pores touch more grains. In that case the shape is close to a rounded polyhedron with vertices located at the enclosing grain triple lines in order to satisfy the surface tension balance; a tetrahedral example is shown for the middle row. This is in line with Tikare2001d , HOTZER2016 , who employed a more regular geometry with pores being placed on quadruple points between a basal hexagonal polycrystal and one top crystal growing into this polycrystal. They observed that the pores took on shapes similar to ice-cream cones, with a triangular cross-section, which is the cross-section of a tetrahedron. Pores merging results in temporary irregular shapes as shown on the bottom. This pore is the result of the tetrahedral pore (four-sided face showing) from the middle row merging with the spheroidal pore shown in the middle panel and one nearby other tetrahedral pore. A video of the time evolution of this subdomain is deposited with the supplementary material. Due to the intentionally large pore mobility no pores are left behind within the growing grains.
@@
For pure grain growth, the speedup can likely be seen as a lower limit for the simplest phase-field problems, as the large amount of evolving interfaces requires fine temporal resolution to solve accurately before transitioning into a more regular behaviour. Adding mobility or interfacial energy contrasts or anisotropy will increase the speedup as the stiffness increases. External coupling to a known temperature ramp to reproduce experimental results would achieve even larger speedups since while the external timescale is fixed, the STS integrator can take dynamics-appropriate timesteps to cover the external timescale.
With particles, the diffusion value enters the stiffness question, but at the same time the error in the concentration evolution will be affected too. Seeing the immobile particles case as the zero diffusion limit of the mobile particles case, this gives us a lower limit for the speedup, with increasingly larger diffusion coefficients increasing the speedup.
6 Conclusion & outlook
@@
The speedups of the STS schemes relative to the FEuler scheme given in the work can typically be seen as lower limits for the respective benchmark cases and the applications which they represent. The benchmarks relating to grain growth can be taken as precursors to sintering, where advective terms need to be added as described above. The Stefan problem is the prototype for phase-changes problems such as dendritic and eutectic solidification in higher dimensions; the speedup reported here should readily transfer to these. As the time scale on which dynamics occurred was not far removed from the time scale induced by the largest eigenvalue of the right-hand side, there was no extra stiffness biasing the results against the FEuler scheme. Only for the case of mobile pores did these time scales differ substantially, which resulted in larger speedups.
What bears further investigation is the relatively large number of step rejections for some benchmarks. This suggests that either the stability limit is passed or that some part (error estimator, time step adapter) in the step adaption machinery fails. The former might be probed by including a dynamical estimate of the largest eigenvalue of the system. If this doesn’t affect the rejection rate significantly, it might be that standard adapters need significant tuning in their coefficients to reduce their rejection rate as shown by Ranocha2022 for compressible CFD problems.